Decide whether the statement is possible or impossible for an angle θ.

cos θ = 5/7 and sec θ = 7/5

A) Possible B) Impossible

I have the correct answer but I need the problem worked out to know why it is possible.
Answer #1

A) Possible

Recall the Quotient formula, sec θ =1/cosθ

So, basically secθ is reciprocal of cosθ .

Given, cos θ = 5/7, hence using sec θ =1/cosθ

sec θ =1/(5/7) =7/5